### v2

# ViLLa: Video Reasoning Segmentation with Large Language Model

Abstract
Although video perception models have made remarkable advancements in recent years, they still heavily rely on explicit text descriptions or pre-defined categories to identify target instances before executing video perception tasks. These models, however, fail to proactively comprehend and reason the user’s intentions via textual input. Even though previous works attempt to investigate solutions to incorporate reasoning with image segmentation, they fail to reason with videos due to the video’s complexity in object motion. To bridge the gap between image and video, in this work, we propose a new video segmentation task — video reasoning segmentation. The task is designed to output tracklets of segmentation masks given a complex input text query. What’s more, to promote research in this unexplored area, we construct a reasoning video segmentation benchmark. Finally, we present ViLLa: Video reasoning segmentation with Large Language Model, which incorporates the language generation capabilities of multimodal Large Language Models (LLMs) while retaining the capabilities of detecting, segmenting, and tracking multiple instances. We use a temporal-aware context aggregation module to incorporate contextual visual cues to text embeddings and propose a video-frame decoder to build temporal correlations across segmentation tokens. Remarkably, our ViLLa demonstrates capability in handling complex reasoning and referring video segmentation. Also, our model shows impressive ability in different temporal understanding benchmarks. Both quantitative and qualitative experiments show our method effectively unlocks new video reasoning segmentation capabilities for multimodal LLMs. The code and dataset will be available at https://github.com/rkzheng99/ViLLa.
1 Introduction
Figure 1: We propose a new video segmentation task that involves complex reasoning for models. Our ViLLa is an effective and efficient LLM capable of segmenting and tracking multiple objects while understanding the user’s implicit instructions.
Capitalizing on the achievements of Large Language Models (LLMs) vicuna ; chatgpt ; openai2023gpt4 ; zhang2023llama , the development of large multimodal models (LMMs) Qwen-VL ; li2023blip ; liu2023llava ; 2023interngpt ; wang2023internvid ; zhu2023minigpt has notably enhanced visual perception capabilities and user-interaction experiences to new heights. However, in the absence of explicit instructions specifying target objects or categories, the majority of models alayrac2022flamingo ; li2023otter ; li2023blip ; liu2023improved ; liu2023visual ; ye2023mplug ; zhu2023minigpt , including both perception models and LMMs, fall short of generating detailed and granular responses such as instance masks, and instead produce merely general textual descriptions for images or videos. This failure, unfortunately, hinders the practical application of multimodal systems in substantial industrial and daily life scenarios, such as autonomous driving, image/video editing, robotics, and augmented reality.
Recent studies lai2023lisa ; ren2023pixellm ; yang2023improved have explored the implementation of LLMs to produce object masks in a novel reasoning segmentation task, which enhances the applicability to real-world applications. Unlike traditional segmentation, which explicitly specifies target objects (e.g., “orange”), reasoning segmentation necessitates complex reasoning to interpret more sophisticated instructions (e.g., “the fruit high in Vitamin-C”). These new approaches propose efficient LMMs that are capable of producing multiple open-set targets while maintaining diverse reasoning complexities. Albeit the success gained by these methods, they still merely focus on image-level reasoning and understanding, which fail to extend the fine-grained reasoning capabilities to the temporal dimension. The adaptation from image to video seems intuitive but still has a major obstacle: The complex relations inherent in video frames pose a challenge for image-based reasoning methods to effectively model the temporal relationships between instances using a single segmentation token or fixed segmentation codebook. Yet, concentrating on modeling temporal relationships will compromise the spatial comprehension per frame. Considering the success of using LLMs in image-level reasoning segmentation, we intuitively come up with the question: how to efficiently utilize the temporal relations across frames while maintaining the frame-level comprehension when we manage to adapt LLMs to video scenarios?
In this paper, we introduce ViLLa, an effective and efficient LMM for video reasoning segmentation and understanding which proficiently perceive and reason videos without losing the spatial details of images (As shown in Fig 1). Our core idea is to represent the videos with both temporally correlated and contextually preserved tokens. Specifically, we propose a context attention module to aggregate frame context information from visual features to text embeddings to generate contextually enriched text embeddings. To incorporate temporal relations to segmentation tokens, we design a video-frame decoder to communicate the segmentation embeddings of a single frame with the embeddings representing the whole video frame. Remarkably, ViLLa demonstrates competitive performance among all video understanding benchmarks, demonstrating the robustness of our models. As shown in  1, our model shows capability in various complicated video scenarios involving complex reasoning and multi-target tracking while segmenting. Tested on Refer-Youtube-VOS and Youtube-VIS, our model could surpass previous state-of-the-art methods by 3.6 and 2.8 in $\mathcal{J}\&\mathcal{F}$ and AP.
In addition, to validate its effectiveness, we establish a benchmark for video reasoning segmentation evaluation, called VideoReasonSeg. Comprising over one thousand video-instruction pairs, this benchmark offers persuasive evaluation metrics for the task. Utilizing a GPT-4V openai2023gpt4 -aided data curation pipeline, we generate 1934 video samples for video reasoning segmentation in either the form of multiple-choice QAs or video-instruction-mask data samples.
In summary, our contributions are as follows:
  • We introduce the reasoning video segmentation task, which necessitates pixel-level video reasoning based on implicit user instructions. Developing such capability is a meaningful leap forward for building an omnipotent video perception model.
  • We have developed VideoReasonSeg, a comprehensive benchmark for video reasoning segmentation that comprises 1934 video-instruction-mask data samples. This benchmark plays a crucial role in evaluating and fostering the exploration of reasoning capabilities in video-based models by the research community.
  • We present our model, ViLLa, which serves as a novel LMM for video reasoning segmentation. It handles tasks with diverse reasoning complexities, achieving state-of-the-art results across various video understanding benchmarks and our proposed benchmark.
2 Related Works
Large Multimodal Models. Large multimodal models (LMMs) have made significant strides in tasks that require understanding diverse modalities. These models can be broadly categorized into two groups based on their utilization of large language models (LLMs).
The first category comprises models lu2022unified ; wang2022ofa ; yu2022coca that are either trained from scratch or leverage smaller language models like BERT for text processing. These models typically employ a combination of contrastive and generative objectives to tackle a range of multimodal tasks (e.g., Coca yu2022coca ). However, their limited language understanding capacity often restricts their performance in tasks that demand massive commonsense reasoning abilities.
The emergence of LLMs in the recent few years has paved the way for a new paradigm in LMM development, in which LLMs are augmented with multimodal comprehension abilities zhang2022opt ; touvron2023llama ; alpaca ; peng2023instruction . This approach typically includes utilizing adapters to align visual and textual representations within LLMs, as exemplified by models such as Flamingo alayrac2022flamingo , BLIP-2 li2023blip , MiniGPT-4 zhu2023minigpt , llama-adapter llama-adapter-v2 , LLaVA liu2023llava , InstructBLIP instructblip , InternGPT 2023interngpt , QwenVL Qwen-VL , InternVideo2 wang2024internvideo2 . Some video-related MLLM (VideoLLM) VideoChatGPT maaz2023video , and Valley luo2023valley utilize ChatGPT chatgpt to generate video instruction-tuning data, aiming to enhance instruction-following capabilities for real-world video comprehension. Even though these models have demonstrated improved performance in vision-language tasks through instructional tuning, their primary limitation lies in generating only textual outputs about the entire image or video, thereby constraining their applicability in tasks that require more fine-grained, region-level, or pixel-level understanding.
Video Segmentation. Video Instance Segmentation (VIS) yang2019video is a fundamental while challenging vision task that aims to detect, segment, and track object instances inside videos based on a set of predefined object categories simultaneously. With the prosperous video media, VIS has attracted various attention due to its numerous important applications in areas such as video understanding, video editing, autonomous driving, etc. Numerous studies athar2020stem ; lin2021video ; wu2022efficient ; wang2021end ; cheng2022masked ; heo2022vita ; wu2022seqformer ; heo2022generalized ; li2022videoknet ; OMGSeg ; dvisdaq have proposed diverse designs (such as tracking by detection, masked attention, object token association, contrastive memory, and more) to effectively model temporal relationships while classifying instances. To name a few, IDOL wu2022defense , which built upon Deformable-DETR zhu2020deformable , incorporates a contrastive learning head to obtain unique embeddings for associations he2020momentum . Subsequently, CTVIS ying2023ctvis improved IDOL by establishing a unified framework for training and inference. DVIS zhang2023dvis , on the other hand, proposed a decoupling strategy for VIS by dividing it into three independent sub-tasks: segmentation, tracking, and refinement.
The Referring Video Object Segmentation (RVOS) task was introduced to segment the target instance referred by a given text in given videos. Early works khoreva2019video ; seo2020urvos used the spatial-temporal memory mechanism to strengthen temporal referring relations. The researchers then started building off-line reference models wang2019asymmetric ; wang2020context ; wu2022multi ; yan2024referred that take the whole clip as input. Currently, query-based Transformer models botach2022end ; wu2022language have begun to dominate. MTTR botach2022end utilized an instance-level segmentation pipeline to predict sequences for all instances and determine the most suitable sequence corresponding to the referred object. ReferFormer wu2022language transformed text expressions into queries to attend to relevant regions in videos.
Reasoning Segmentation. In numerous practical applications, it is essential to comprehend visual inputs at a more fine-grained level, such as specific regions or even individual pixels. To address this need, several approaches have been proposed to equip Large Language Models (LLMs) with fine-grained understanding capabilities. Kosmos-2 peng2023kosmos , InternGPT 2023interngpt , and Ferret you2023ferret offer grounding capabilities to specified image regions, typically encoding location coordinates as tokens for integration with LLMs. Yet, these approaches still fail to provide pixel-level outputs. To bridge this gap, LISA lai2023lisa integrates SAM kirillov2023segment with LLMs to tackle segmentation tasks. Furthermore, LISA pioneered the use of Large Multimodal Models (LMMs) for complex instruction reasoning, a paradigm shift from traditional tasks that rely on explicit human instructions for object or category identification. This innovative approach enables LMMs to reason about intricate instructions, moving beyond simplistic object recognition. In order to generate masks for reasoning tasks involving multiple targets, PixelLM ren2023pixellm proposes a novel pixel decoder and a comprehensive segmentation codebook to conduct multi-target reasoning segmentation. However, these methods mainly focus on reasoning image segmentation and fail to reason and deduct at a pixel-level across temporal dimensions. When they are transferred to the video domain, the segmentation token is not capable of modeling multiple targets as well as their moving trajectories. To bridge this gap, we propose ViLLa with context aggregation and video-frame decoder to address the problem of video reasoning segmentation.
3 Video Reasoning Segmentation
3.1 Problem Definition
Video reasoning segmentation task aims to output a binary segmentation mask sequence $\{m_{t}\}_{t=1}^{T}\in\mathbb{R}^{T\times H\times W}$, given an input video clip consisting of $T$ frames $V\in\mathbb{R}^{T\times H\times W\times 3}$ and an implicit query text instruction $\mathbf{x}_{txt}$. The task shares a similar formulation with the referring video object segmentation, but there’s a major difference that makes video reasoning segmentation much more challenging: The complexity of the input query text. Apart from the straightforward text description, such as “a zebra to the left”, the query text of video reasoning segmentation includes more complex and nuanced descriptions of the instance (e.g., “a type of African mammal known for their distinctive black and white striped coat patterns walking at the end of its herd”), which involves complicated reasoning as well as the world knowledge. Also different from the image reasoning segmentation, the query text of video reasoning segmentation includes more motion information, such as “a person wearing a white shirt is in the ocean on a surfboard riding a wave”, which requires the model to capture the moving trajectory of the “person riding a wave”. Thus, this topic is challenging and worth exploring, especially in a time when video inputs have become ubiquitous.
3.2 Benchmark
Given the lack of quantitative evaluation, it is necessary to establish a benchmark for the video reasoning segmentation task. To ensure reliable assessment, we have collected a diverse set of videos from Youtube-VIS yang2019video , OVIS qi2022occluded , and LV-VIS wang2023LVVIS , annotating them with implicit text instructions and high-quality target masks. To fully evaluate the reasoning capabilities, we design two types of evaluation standards: 1) multiple-choice QAs; 2) instructions and answers. Considering that not all the annotations of selected datasets follow the multiple-choice QA format, we automatically convert the video annotations into this format via LLMs. This is mainly because the open-ended answer has to be scored by LLMs or user studies, which may either introduce evaluation bias or manual intervention. Ultimately, we produce 2 multiple-choice QA pairs for each of the video clips. As for the answers, we mainly focus on the quality of segmentation masks. The resulting VideoReasonSeg benchmark comprises a total of 1934 video-instruction-mask data samples. This dataset is further partitioned into three splits: train, val, and test, containing 1000, 400, and 534 video samples, respectively. The details of data annotation are given in the supplementary material.
Dataset Generation Pipeline. Inspired by PixelLM, we use GPT-4V and Qwen-VL for generating questions and multiple choices, because of their superiority in understanding visual contents. These models have been instrumental in generating more nuanced, contextual, and naturalistic questions. Specifically, we feed all the instance category names and corresponding masks in the video information in the video to GPT-4V. With our carefully designed prompts, GPT-4V autonomously selects instances to construct question-answer pairs based on the video contents. Examples of such prompts are illustrated in the supplementary files.
4 Method
Figure 2: Overall framework of the proposed ViLLa. Given the input video frames and user input text query, the context attention module firstly aggregates text-related visual cues from the visual features to the text embeddings. The large language model generates text output and segmentation tokens with the input of visual features, visually enriched text embeddings, and pre-defined segmentation tokens. Finally, the segmentation tokens are fed to the Video-Frame Decoder to produce the final output segmentation mask tracklets.
4.1 Model Design
Model Overview. As depicted in Fig. 2, ViLLa features a streamlined architecture, comprising mainly several main parts: 1) a pre-trained visual encoder that aligns with a text encoder 2) a large language model 3) a context attention module and 4) a video-frame interaction decoder.
While 1) and 2) components follow well-established large language model (LLM) architectures to ensure compatibility, the context aggregation module (CAM) and the Video-Frame decoder are crucial in equipping the LLMs with the capacity to generate masks across temporal dimensions. We utilize the context aggregation module to aggregate the target-text-relevant visual features of the current frame to the input user text embeddings. The output text embeddings that carry the most crucial visual cues are then fed into the large language models along with the visual features. In video-frame decoder, we use the video-scale segmentation tokens to interact with frame-level tokens to produce the final refined complete segmentation tokens that contain both video-level and frame-level information. Finally, the segmentation tokens produce the prediction masks in conjunction with the multi-scale video features.
Encoders. For input frame $V_{\text{t}}$, the visual encoder $\mathcal{I}$ extracts multi-scale visual features $F_{{t}}=\{F_{{t}}^{\ell}\}_{\ell=1}^{L}$ from $\mathcal{I}(V_{{t}})$, comprising $L$ visual features output at select layers of $\mathcal{I}$. Meanwhile, the user instruction is fed into the text encoder with the visual embedding $F_{t}^{L}\in\mathbb{R}^{N\times C}$ (where $N=H/p\times W/p$ and $C$ indicate the number of frame patches and embedding channels, respectively) and generates the $\mathbf{x}_{txt}\in\mathbb{R}^{M\times C}$, where $M$ denotes the number of queries. In this part, we choose QFormer as the text encoder to maintain the cross-modality alignment and interaction. By incorporating this approach, the text embeddings now capture visual cues that are highly relevant to the user’s instructions. The output of the final layer, $F_{t}^{L}$, encodes global image information and is transformed to align with the language space of LLMs via a vision-to-language projection layer $\text{Proj}_{V\rightarrow L}$.
Context Aggregation Module. In this module, we aim to aggregate text-related visual features and inject them to generate text embedding that could represent the current frame. With the text embeddings $\mathbf{x}_{txt}$ and visual features, we formulate the context-based aggregation as:
$\textbf{E}_{t}=\text{FFN}(\text{CrossAttn}(\mathbf{x}_{txt},F_{t}^{L},F_{t}^{L})),$ (1)
$\textbf{E}_{t}^{c}=\text{Concat}\{\textbf{E}_{t}^{i}\}_{i=1}^{K},$ (2)
where ‘CrossAttn’ refers to cross-attention operation, and $\mathbf{x}_{txt}$ is the query, while $F_{t}^{L}$ is the value and key. By conducting the cross-attention, we further incorporate the visual cues into the refined text embeddings $\textbf{E}_{t}=\text{Concat}\{\textbf{E}_{t}^{i}\}_{i=1}^{M}$. However, we argue that not all queries are needed for the refined text embeddings. Unlike QFormer which adopts 32 input queries as input LLM tokens, we propose to condense these embeddings into more concentrated tokens. After getting $\textbf{E}_{t}$ from the context aggregation between text embeddings and visual features, we select the output embeddings with $K$ higher-response scores from the attention matrix. That is, the final condensed input embeddings $\textbf{E}_{t}^{c}$ preserve the most relevant visual information related to user input.
Multi-level Segmentation Tokens. Aiming to enrich the encoding of target-specific information and thereby facilitate the generation of high-quality mask tracklets across temporal dimensions, we devise multi-scale segmentation tokens, representing both frame-level and video-level concepts, tailored to meet the characteristics of video segmentation tasks that demand the modeling of multiple target movements. Formally, we define $C_{\text{seg}}=\left\{c_{n}^{s}\in\mathbb{R}^{d}\right\}_{n=1}^{N}$, where $s\in\{f,v\}$ indicates the frame- or video-scale tokens, $N$ denotes the number of tokens per scale, and $d$ refers to the hidden dimension of LMMs. The multi-level segmentation tokens $C_{\text{seg}}$, combined with the visual features $F_{t}^{L}$, and condensed text embeddings $\textbf{E}_{t}^{c}$, are then processed by the LLM to autoregressively generate the response $y_{\text{out}}$:
$y_{\text{out}}=\mathcal{F}(\text{Proj}_{V\rightarrow L}(F_{t}^{L}),\textbf{E}_{t}^{c},C_{\text{seg}}).$ (3)
Using an example can provide a more vivid explanation of how Large Language Models generate responses. When the user inputs “segment the leftmost zebra in the video”, the output will include not only a textual answer but also the segmentation tokens. $C_{\text{seg}}$: “The zebra is $c_{1}^{f},c_{1}^{v}$” (considering the easiest case where $N=1$ for $C_{\text{seg}}$). Before the decoder, an additional projection layer $\Phi$ is utilized to adjust the dimensions of segmentation tokens, in the form of $Q^{s}=\Phi(C_{\text{seg}})$, ($s\in\{f,v\}$). The output multi-scale segmentation embeddings $Q^{v},Q^{f}$ derived from $C_{\text{seg}}$ are inputs to the pixel decoder $\mathcal{D}$ alongside visual features $F_{{t}}$ for mask tracklet generation.
Video-Frame Decoder. We design a novel Video-Frame decoder $\mathcal{D}$ to generate mask tracklets from the multi-scale visual features $F_{{t}}$ and segmentation tokens $Q^{v},Q^{f}$. Based on the design of the transformer decoder in Mask2Former, $\mathcal{D}$ consists of $L$ decoder layers, where $l^{\text{th}}$ layer cascades a masked cross-attention $h_{\text{CA}}^{l}$, a self-attention $h_{\text{SA}}^{l}$, and a feed-forward network $\text{FFN}^{l}$. Now, we have the two scales of segmentation embeddings, $Q^{f}$ focusing on every frame separately, while $Q^{v}$ interacting with the whole video features.
Thus both the frame- and video-level segmentation embeddings pass the transformer decoder. For each step, these two embeddings are updated as follows:
$\mathbf{Q}^{l+1,s}=\text{FFN}^{\text{s}}(h_{\text{SA}}^{\text{s}}(h_{\text{CA}}^{\text{s}}(\mathbf{Q}^{l,s},{F_{t}^{l}}))),$ (4)
where $s\in\{f,v\}$, and the $h_{\text{CA}}^{s}(q,r)$ indicates the cross-attention with query embedding $q$ and reference embedding $r$. Then, we feed the frame- and video-level embeddings in aggregation in order to interchange video-level and frame-level information, and this process can be formulated as:
$\mathbf{Q}^{l+1,v^{{}^{\prime}}}=\gamma\cdot\text{Softmax}(({Q}^{l+1,v}\times{{Q}^{l+1,f}}^{\top})\times{{Q}^{l+1,f}})+(1-\gamma)\cdot\mathbf{Q}^{l+1,v},$ (5)
where $\mathbf{Q}^{l+1,v^{{}^{\prime}}}$ is the modulated video-level segmentation embeddings at $l^{\text{th}}$ layer. In this aggregation process, we aggregate the frame-level embeddings with higher responses to our video-level segmentation embeddings in a momentum-based manner. The momentum factor $\gamma$, which is empirically set to 0.03, governs the update of the video-level embeddings. This choice is based on the assumption that the aggregation of frame-level embeddings should not induce substantial changes to the overall video-level representations.
Training Objectives. The model is trained end-to-end using the text generation loss $\mathcal{L}_{txt}$ and the segmentation mask loss $\mathcal{L}_{mask}$. The overall objective $\mathcal{L}$ is the weighted sum of these losses, determined by $\lambda_{txt}$ and $\lambda_{mask}$:
$\mathcal{L}=\lambda_{txt}\mathcal{L}_{txt}+\lambda_{mask}\mathcal{L}_{mask}.$ (6)
where mask loss $\lambda_{mask}$ can be further divided into binary cross-entropy loss and the dice loss for both video-level and frame-level: $\mathcal{L}_{mask}^{s}=\lambda_{ce}^{s}\mathcal{L}_{ce}^{s}+\lambda_{dice}\mathcal{L}_{dice}$ ($s\in\{f,v\}$), and $\mathcal{L}_{txt}$ is the auto-regressive cross-entropy loss for text generation.
5 Experiment
In this section, we first present the implementation details and the evaluation benchmarks as well as baselines, and then show the comparison results on different video perception and our proposed reasoning benchmarks. Finally, we ablate on the key components in ViLLa.
5.1 Implementation Details
We use pre-trained multimodal model from InternVideo2-1B wang2024internvideo2 and InternVideo2-6B, with LoRA adopted for efficient fine-tuning. For the visual encoder and LLM, we apply pre-trained UMT-L from the InternVideo2 stage1 model as the vision encoder and Vicuna-7B v0. Following BLIP2, we deploy QFormer using the pre-trained $BERT_{base}$. To generate multi-scale visual features, we adopt the ViT-Adapter to generate the necessary multi-scale features for the video-frame decoder. It is important to note that video segmentation differs from image segmentation in that video segmentation involves capturing inter-frame relationships across multiple frames during training. This added complexity results in higher GPU memory requirements compared to image segmentation. To address this challenge, an efficient version of the VIT-Adapter is applied, which helps to reduce the memory consumption of the video segmentation model by removing all injectors.
We use 8 A100 GPUs for training 50 epochs (approximately 1.5 days). We use the AdamW optimizer with the learning rate and weight decay set to $52e-$ 5 and $0.02$, respectively. The batch size per device is set to 4, the input frame number is 8, the input resolution is 224, and the gradient accumulation step is set to 10. Further details are presented in appendix.
5.2 Benchmarks and Baselines
Benchmarks. We evaluate ViLLA on benchmarks with video segmentation and question-answering evaluations: VideoReasonSeg, referring video segmentation (Refer-Youtube-VOS, Ref-DAVIS17), conventional video segmentation (Youtube-VIS series), and VQA datasets in supplementary files (FunQA, Perception Test, CLEVRER). Through this evaluation, we validate the versatility of ViLLa in diverse segmentation tasks and question-answering tasks. For the VideoReasonSeg benchmark, we use the average precision (AP) defined in video segmentation as well as multiple choices to evaluate the video reasoning capabilities of models. In video segmentation, we formulate the queries following the dataset’s annotations in the format of “Please segment the <description>”, with <description> corresponding to the descriptions of the instance categories. In referring video segmentation, we follow previous methods to measure the $\mathcal{J}\&\mathcal{F}$. Training involves random sampling from VideoReasonSeg, VQA data, Video instance segmentation, and Referring video segmentation datasets. The model also follows the training setting of PixelLM ren2023pixellm to pre-train on image datasets to provide an initial value for segmentation tokens.
Baselines. To our best knowledge, ViLLa is the first to handle video reasoning segmentation. To demonstrate the effectiveness of ViLLa, we establish strong baselines for comparative analysis on the mentioned benchmarks. These baselines are developed from previous reasoning segmentation approaches (LISA, PixelLM), referring video segmentation methods (ReferFormer, MTTR), and video segmentation models (VITA, Mask2Former).
Table 1: Reasoning video segmentation results among ViLLa and previous related works on VideoReasonSeg benchmark. “Segm” refers to ‘Segmentation’ while “MC” indicates “Multiple Choices”.
Method Val Test
Segm MC Segm MC
AP AR Accuracy AP AR Accuracy
Mask2Former cheng2022masked 37.9 33.2 35.8 36.5 32.8 34.3
VITA heo2022vita 39.3 34.9 36.5 38.8 34.1 35.7
MTTR botach2022end 38.3 32.9 36.4 36.9 32.7 35.0
ReferFormer wu2022language 39.5 34.4 37.0 39.0 33.9 35.9
LISA-Llama-7B lai2023lisa 18.3 14.9 34.5 16.7 12.9 33.0
LISA-Llama2-13B lai2023lisa 20.1 15.4 35.2 18.5 13.6 34.7
PixelLM-Llama2-7B ren2023pixellm 33.2 29.5 42.5 30.4 29.7 37.2
PixelLM-Llama2-13B ren2023pixellm 34.2 30.1 44.6 31.8 29.9 41.3
ViLLa-InternVideo2-1B wang2024internvideo2 42.5 35.5 50.1 39.8 37.4 47.8
ViLLa-InternVideo2-6B wang2024internvideo2 44.3 39.2 53.4 41.2 38.8 50.5
5.3 Main Results
Video Reasoning Segmentation. Tab. 1 compares ViLLa with adapted competing methods on our video reasoning segmentation task. For the image reasoning segmentation method, to adapt to the video segmentation scenario, we add an extra tracker based on SAM-Track cheng2023segment . Since the original LISA only has one segmentation token, we calculate the AP between the predicted one tracklet and the ground truth tracklets. As for the VIS methods, we add the LLM to maintain their capability of complex reasoning.
As shown in Tab. 1, our method outperforms the previous image reasoning segmentation approaches by a very large margin, demonstrating that our model is capable of accomplishing tasks involving complex reasoning across time. Because video reasoning, different from referring video segmentation and image reasoning segmentation, requires the model to possess reasoning ability based on the whole video. Thus, without comprehensive knowledge of the whole video, the model can not perform well. Also, we compared our method with two-stage methods which use LLM to generate text while using a video segmentation model to produce the tracklet. The results also indicate that our method exceeds these two-stage methods significantly. We argue that this is because we are trained end-to-end, while these methods decouple the video modeling and reasoning. Also, our model improves the association between video and frame via our specially designed context aggregation strategy and decoder structures, which distinguishes our method from previous works.
Table 2: Comparison on Ref-YouTube-VOS and Ref-DAVIS17.
Method Ref-YouTube-VOS Ref-DAVIS17
[…]
ReferFormer wu2022language 62.9 61.3 64.6 61.1 58.1 64.1
OnlineRefer wu2023onlinerefer 62.9 61.0 64.7 62.4 59.1 65.6
LISA lai2023lisa 63.8 62.1 65.6 62.5 59.2 65.8
PixelLM ren2023pixellm 65.0 63.3 66.8 63.1 60.4 66.0
ViLLa 66.5 64.6 68.6 64.4 61.2 67.7
Table 3: Comparison on VIS datasets (“YTVIS” refers to “Youtube-VIS”).
Method YTVIS-19 YTVIS-21 OVIS
AP AP AP
SeqFormer wu2022seqformer 59.3 51.8 -
Mask2Former cheng2021mask2former 61.6 55.3 24.1
VITA heo2022vita 63.0 57.5 27.7
IDOL wu2022defense 64.3 56.1 42.6
ViLLa 67.1 59.5 46.2
Referring Video Segmentation. Tab. 3 presents the results on the referring video segmentation datasets. Refer-YouTube-VOS seo2020urvos expands Youtube-VOS xu2018youtube using textual descriptions. It consists of 3,975 videos and 27,899 expressions. Refer-DAVIS khoreva2019video extends DAVIS17 pont20172017 , which has 90 videos and more than 1,500 expressions. As shown in the table, our ViLLa demonstrates better performances on these two datasets, surpassing previous referring segmentation models and image-reasoning methods by 3.6 and 1.5 in $\mathcal{J}\&\mathcal{F}$ on Ref-Youtube-VOS dataset, respectively.
Video Instance Segmentation. Tab. 3 presents the results on the video instance segmentation datasets. YouTube-VIS 2019 yang2019video , contains 2.9k videos. The dataset was updated to YouTube-VIS 2021 with longer videos. OVIS dataset is another resource for video instance segmentation, particularly focusing on scenarios with severe occlusions between objects qi2022occluded . It consists of 25 object categories and 607 training videos. Our ViLLa surpasses previous SOTA VIS methods by 2.8, 3.4, and 3.6 points, respectively. The results prove that our model is excelling at modeling temporal relations and segmenting high-quality tracklets.
Table 4: Ablation study on the design choice of vision backbone.
Vision Backbone Segmentation Multiple Choices
AP AR Accuracy
SAM kirillov2023segment 40.5 36.4 46.0
Mask2Former-SwinL cheng2022masked 39.6 35.3 45.7
InternViT  chen2024far 43.5 38.7 52.9
InternVideo2 wang2024internvideo2 44.3 39.2 53.4
Table 5: Ablation study on key component designs in ViLLa.
ID Context Video-Frame VidReasonSeg Refer-VOS
Aggregation Decoder AP Accuracy $\mathcal{J}$$&\mathcal{F}$
1 40.6 49.5 63.7
2 41.5 50.8 64.2
3 44.3 53.4 66.5
Table 6: Ablation study on key designs of Video-Frame Decoder.
ID Multi-scale Video-frame VidReasonSeg Refer-VOS
Tokens Aggregation AP Accuracy $\mathcal{J}$$&\mathcal{F}$
1 41.5 50.8 64.2
2 42.4 50.6 65.0
3 44.3 53.4 66.5
Table 7: Ablation study on the scales.
Architecture VidReasonSeg Refer-VOS
AP Accuracy $\mathcal{J}$$&\mathcal{F}$
Baseline 41.5 50.8 64.2
+ 2 scale frame-video aggregation 43.8 52.5 66.1
+ 3 scale frame-video aggregation 44.3 53.4 66.5
+ 4 scale frame-video aggregation 44.2 53.0 66.4
5.4 Ablation Studies
In this section, we conduct extensive ablation studies to reveal the contribution of each component. Unless otherwise specified, we conduct the experiments on our VideoReasonSeg dataset.
Design Choices of Backbone. We emphasize that different vision backbones can be applicable in our framework. In Tab. 5, we show that Internvideo2 performs the best, potentially because it is the state-of-the-art video foundation model, demonstrating its capacity in multimodal video understanding. Further, with the Mask2Former backbone, our framework still achieves a decent performance on the video reasoning segmentation dataset. This reveals that the design choice of the vision backbone is flexible and not limited to the video foundation model.
Key Components Design. Tab. 5 demonstrates the effect of our component designs based on the vision backbone and transformer decoder in Mask2Former. By adopting our algorithm, we significantly improve the baseline model by strengthening the video-frame relationship and fusing context information into text embedding. Tested on VideoReasonSeg, our context aggregation improves the accuracy of answering multiple choices by 1.3 points, which indicates that adding context information to the text embedding and keeping the textual input contextualized is beneficial for video understanding. Also, the design of video-frame decoder improves the segmentation performance by 2.7 AP on segmentation, showing that improving the video modeling by applying multi-scale tokens and strengthening their relations is crucial.
Video-Frame Decoder. Tab. 7 shows the influence of utilizing multi-scale tokens and video-frame aggregation in video-frame decoder. With multi-scale tokens only, the baseline model achieves a 0.9 AP increase, while applying the aggregation brings about 1.9 AP. This demonstrates that using multi-scale tokens and associating them are beneficial for video segmentation tasks. In Tab. 7, we explore the effects of increasing the number of scales in the decoder. Initially adding layers will contribute to notable gains, but the gain would diminish with more scales of decoder layers.
Figure 3: Comparison between ViLLa, LISA, and PixelLM.
5.5 Qualitative Results
In Fig. 3, we provide a visual comparison with existing related works. We show that our ViLLa is capable of segmenting multiple instances (helmets of riders) while tracking them. On the other hand, LISA and PixelLM, though aggregated with the tracker, can not either segment multiple instances or track all instances properly (as shown in red circles, where they both track the wrong helmet). These qualitative results demonstrate the reasoning and segmentation capacity of our ViLLa.
6 Conclusion
In this study, we propose a new task, video reasoning segmentation. Also, we have introduced a benchmark VideoReasonSeg, which comprises approximately 2k samples. Finally, we introduce ViLLa, an effective and efficient model that aims to provide some early explorations towards this task. Benefiting from our exquisite designs, ViLLa demonstrates convincing capabilities in producing temporally consistent high-quality masks for a variety of video segmentation tasks.
Limitations and broader impacts. Even though our model performs well in reasoning segmentation with text instructions, the user input is still limited to text, rather than a variety, such as scribbles. Also, our model can not support multi-round conversations. However, We hope our work can provide new insights into the future direction of combining LLMs and vision tasks so as to build an omnipotent perception model.
Acknowledgement. This work is partially supported by the National Natural Science Foundation of China (No. 62201484), National Key R&D Program of China (No. 2022ZD0160100), HKU Startup Fund, Shanghai Artificial Intelligence Laboratory, and HKU Seed Fund for Basic Research.
Appendix
This appendix provides more details about the proposed ViLLa, and our proposed benchmark, VideoReasonSeg. The first part includes data details while the second part contains the data generation pipeline of our VideoReasonSeg dataset, and in the final part we provide extra ablation experiment.
Appendix A Implementation Details
Training details. In the first part, we present the detailed training configuration in our Tab. 8. As for the $\lambda_{txt}$ and $\lambda_{mask}$, they are set to 1.0, and $\lambda_{dice}$ is 0.5 while $\lambda_{ce}$ is 2.0.
Table 8: Training Hyperparameters for ViLLa.
Config Value
input frame 8
input resolution 224
max text length 512
optimizer AdamW
optimizer momentum $\beta_{1}$, $\beta_{2}$ = 0.9, 0.999
weight decay 0.02
learning rate schedule cosine decay
learning rate 2e-5
batch size 32
warmup iters 10
Appendix B VideoReasonSeg Details
Data Generation Pipeline.
In order to generate multiple-choice QA, we automatically convert the video annotations into this format via LLMs. Specifically, we first use ChatGPT [30] to generate a question for each video. For most questions, we construct the option candidates directly from the ground truth annotations. For example, video segmentation tasks contain masks and instance categories of each video. Then the candidate option for multiple-choices would be the correct category, wrong category, and a not-sure choice. Ultimately, we produce 2 pairs for each of the video. To strengthen the evaluation’s robustness, for each question we randomly sample 3 to 5 answer options from the available candidates and shuffle the order of the options. Additionally, to prevent the common issue of answer leakage where longer options tend to be correct, we further use a large language model to ensure that all the answer options for a question are of similar and reasonable lengths.
As for the question and answer pair, we use GPT-4V and Qwen-VL to construct our dataset. We utilize videos with pre-existing video mask annotations. The video frames, the category names contained in the video, and their related mask annotation are contained in the prompts feed to the GPT-4V. An example of the prompts is shown in Fig. 4. Using carefully crafted prompts, GPT-4V autonomously selects instances to construct question-answer pairs relevant to the video. As illustrated in Fig. 5, we demonstrate two types of question, both question and multiple choice, from a given video. In this example, we show that our data tests the capability of models of reasoning based on common world-knowledge, and relate ‘vehicle carrying passengers’ to the white bus on the roadside. In addition, the multiple choice expects the model to distinguish the type of vehicle from other plausible answers, such as ‘taxi’ and ‘bike’. All together these questions are tests of the reasoning capacities of models on both pixel-level and video-level.
Although GPT-4V can efficiently understand the content of the video frame, there are still failure cases in the generated data. One major problem is that questions can be too objective and hard to evaluate. For example, the question “How would you rate the overall difficulty and impressiveness of the skateboarding you observed?” is very objective and the answers can be various for different people. And this requires further prompts and filtering during the data generation process.
Visualizations of ViLLa. Extra visualizations are presented in Fig. 6, which show a case that successfully segments a brown cow moving forward.
Figure 4: GPT-4V data generation pipeline. The right part shows an example of how reasoning segmentation data and multiple choices are generated. The input prompt includes certain rules and the position as well as time localizations to instruct GPT-4V into generating more effective data samples.
Figure 5: GPT-4V data generation samples. The part shows further samples of the generated questions and the multiple choices. The two types of questions can help we better evaluate the model’s performance in video reasoning from both pixel-level and video-level.
Figure 6: Visualizations of ViLLa.
Appendix C Ablation Studies
Aggregation Strategy. Tab. 9 shows the results of different aggregation strategies in the video-frame decoder. We compare our aggregation between video-level segmentation embeddings with the feature fusion adopted in PixelLM [36]. As shown in the table, our strategy improves the performance on referring VOS dataset, demonstrating the effectiveness of the video-frame aggregation strategy.
Table 9: Ablation study on aggregation strategies.
Strategy Ref-YouTube-VOS Ref-DAVIS17
[…]
Feature Fusion 65.9 64.1 68.5 63.2 60.5 66.1
Embedding Similarity 66.5 64.6 68.6 64.4 61.2 67.7